OpenStudy (anonymous):

What is the square of 15? Justify the answer please.

OpenStudy (anonymous):

\[15 ^2 = 15*15\]

OpenStudy (anonymous):

In general square of any number a \[a^2 = a*a\]

OpenStudy (anonymous):

Even the square of 5 x 11? 55 x 55

OpenStudy (asnaseer):

This site might be useful to you in understanding squares (and their opposites - square roots): http://www.mathsisfun.com/square-root.html

OpenStudy (anonymous):

@michu \[(5*11)^2= (5*11)(5*11)\]

OpenStudy (anonymous):

so basically you are right @michu